# Topology-Aware Conformal Prediction for Stream Networks

### v1

@@ 6.1 Simulation
Our numerical results demonstrate that our method enhance the predictive efficiency significantly without sacrificing the coverage guarantee, by considering both sample-based and topology-based covariance. From Figure 2, we observe that CP methods employing ellipsoidal uncertainty sets tend to cluster towards the upper region of the plot, indicating higher efficiency compared to CP methods based on spherical or square uncertainty sets. Although MultiDimSPCI achieves the lowest inefficiency, its coverage drops significantly below the required threshold, highlighting its instability when relying solely on the sample covariance matrix. This issue persists even in simulated data designed to align with its error process assumptions. In contrast, with $\lambda$ fixed at $0.5$, our method STACI is positioned near the upper-right corner alongside GT, which leverages the ground-truth covariance matrix. This suggests that our method achieves performance comparable to GT, balancing low inefficiency while maintaining the necessary coverage guarantees. Among all methods that surpass the coverage threshold, our method, STACI($\gamma=0.01$), demonstrates the best efficiency with the smallest variance, further reinforcing its robustness and effectiveness.
Figure 3 reveals a clear trade-off trend between coverage and efficiency: the higher $\lambda$, the confidence level rises, but efficiency declines. This suggests that $\lambda$ must carefully chosen: if too large, our method over-relies on topology and fails to adapt to covariance shift; if too small, it depends more on sample covariance matrices, which are purely data-driven and thus unstable, leading to a coverage drop. Nonetheless, no matter whether adapting confidence level, setting a larger $\lambda$ in STACI can efficiently increase coverage and maintain it near the pre-determined level, while only slightly reducing efficiency, which remains comparable to GT.
Figure 3: Trade-off between coverage and efficiency on synthetic data, where the higher the better performance. The right $y$-axis represents the predictive efficiency, and the efficiency of STACI is accompanied with standard deviation bands.
6.2 Real Data Study
We further conduct experiments on a real-world traffic dataset, Performance Measure System (PeMS) [4], which contains the data collected from the California highway network, providing $5$-minute interval traffic flow counts by multiple sensors, alongwith flow directions and distances between sensors. To model it into stream network, we also rely on [2] to check accurate road connection information. We select $12$ sensors in a South-to-North freeway, and plot their locations and corresponding road segments in Figure 4.
Experiment Configuration
We adopt Adaptive Graph Convolutional Recurrent Network (AGCRN) [1] as the backbone model $f$. We set our default $\lambda=0.5$. For simplicity, we only use fixed weights with all equal values, without requiring any additional information. Multiple hyperparamter and ablation study are also provided over the key parameters in our framework: ($i$) $\lambda$ from 00 to $1$ with step of $0.02$; ($ii$) $n=100,200,300,400,500$; ($iii$) $\gamma=0$ or $0.01$.
Figure 4: Real-world road network structure and its abstraction. The left map displays the road network, where freeways are bold gray lines in blue shade, and ramps off the freeway are represented by blue squares. Based on these ramps and road junctions, the network is divided into different segments. Traffic flow monitoring sensors from $\ell_{1}$ to $\ell_{12}$ are placed exclusively on those northbound freeways, marked with green transparent triangles. The right map provides an abstract representation of the road network and sensor locations, using the same symbols for consistency.
Figure 5: Comparison of methods on PeMS Dataset over coverage ($x$-axis) and efficiency ($y$-axis). The setting is the same as Figure 2, except that the circle size now represents variance of efficiency.
Result
Using 500 calibration samples, Figure 5 shows that among all methods surpassing predetermined confidence level, our method significantly improve efficiency.
As shown in the first line in Figure 6, our methods can greatly alleviate the undercoverage issue, while improving the inefficiency. When $\lambda=0$, the method reduces to using only the sample covariance matrix, which serves as our most competitive baseline, MultiDimSPCI. Our confidence levels are higher than MultiDimSPCI with arbitrary hyperparameters, while the efficiency can also be improved with proper weights. Specially, regardless of calibration sample size $n$, selecting a $\lambda$ from $0.3$ to $0.9$ can always bring better efficiency and coverage, proving the importance of topology information and the robustness of STACI to hyper-parameters.
Another set of ablation experiment results without ACI ($\gamma=0$) is provided in the second line, where similar conclusions can be drawn. Our method can obviously lift confidence level from under $87$% to surpass the desire $95$% level, even with a small CP calibration sample size of 100. From another perspective, when faced with inherent covariance shift over time, incorporating topology information is a robust solution to overcome under-coverage issue and keep informative predictions.
Additionally, we conduct experiments in an offline setting, as presented in Figure 8 of Appendix C.2. In this setting, topology-based and sample-based covariance matrix are not updated when new observations $Y_{t}$ are obtained at the next time $t+1$. Our methods still consistently outperform the strongest baseline, MultiDimSPCI, across a range of hyperparameter selections.
@@ 6.2 Real Data Study
In conclusion, the estimate of the covariance matrix can benefit from both topology and samples, compared with relying on any single resource. The incorporation of both estimators is crucial. First, with limited finite calibration sample $n$, the topology-based estimator offers a more stable structure as it possess fewer parameters. It can alleviate the temporal distribution shift and the resulting under-cover problem, which is also consistent with the theoretical analysis. Second, the sample covariance matrix generally exhibits better efficiency as it more effectively captures the spatial structure of the specific samples in the calibration dataset. However, in many real-world datasets, this approach fails to maintain coverage when there is a temporal distribution shift. This limitation can be mitigated by incorporating a topology-based matrix and adaptively adjusting the significance level.
@@
Figure 6: Comparison of Coverage and Efficiency for PeMS data with different belief weight $\lambda$ and calibration set size $n$, with adaptive step size $\gamma=0.01$ (upper) and 00 (lower). The pre-determined coverage threshold of $95$% is shown by a horizontal gray dotted line.
@@ 7 Conclusion
In this work, we proposed STACI, an adaptive conformal prediction framework for stream networks. Theoretically, we established coverage guarantees and demonstrated the model’s ability to minimize inefficiency under mild conditions. Empirically, STACI produced smaller prediction sets while maintaining valid coverage across both (stationary) simulated data and (non-stationary) real-world traffic data.
There are two potential research directions in the future. First, the STACI framework could be extended beyond stream networks to more general spatio-temporal graphs by replacing $\Sigma_{\mathcal{G}}$ with alternative network parameterizations. This extension would enable the development of novel methods that effectively exploit topological structures in broader spatio-temporal settings. Second, there is room for more in-depth theoretical exploration. Our current analysis mainly ensures approximate average coverage when adaptively calibrating the significance level. Stronger validity guarantees, with explicit finite-sample coverage bounds, could be established by incorporating assumptions on the distribution shift of the error process, such as first-order differencing stationarity.
Appendix A Proof
A.1 Proof of Theorem 5.4
@@ Appendix A Proof
$|\hat{s}_{t}-s_{t}|=|\epsilon_{t}^{\top}A\epsilon_{t}-\hat{\epsilon_{t}}^{\top}\hat{A}\hat{\epsilon_{t}}|\leq|\epsilon_{t}^{\top}A\epsilon_{t}-\epsilon_{t}^{\top}\hat{A}\epsilon_{t}|+|\epsilon_{t}^{\top}\hat{A}\epsilon_{t}-\hat{\epsilon_{t}}^{\top}\hat{A}\hat{\epsilon_{t}}|$
@@
$\leq$ $r\|\Delta\|^{2}+2r\|\Delta\|\|\epsilon\|+o(g(n))\|\epsilon\|^{2}.$ (A.2)
The inequality (i) exists because of the cauchy-schwartz inequality and Assumption 5.2. Hence, by Assumption 5.1, we have
@@
$\leq$ $rn\nu_{n}^{2}+2r\sqrt{n\nu_{n}^{2}(\sum_{t=T-n+1}^{T}\|\epsilon_{t}\|^{2})}+o(g(n))\sum_{t=T-n+1}^{T}\|\epsilon_{t}\|^{2}.$ (A.3)
From Assumption 5.3, we have
$\mathbb{E}\left[\frac{1}{n}\sum_{t=T-n+1}^{T}\|\epsilon_{t}\|^{2}\right]=\frac{1}{n}\sum_{t=T-n+1}^{T}\mathbb{E}[\|\epsilon_{t}\|^{2}]\leq\kappa_{1}I.$ (A.4)
@@
$\mathbb{P}\left(\frac{1}{n}\sum_{t=T-n+1}^{T}\|\epsilon_{t}\|^{2}-\mathbb{E}[\|\epsilon_{t}\|^{2}]\geq\sqrt{\frac{\,\mathrm{Var}[\|\epsilon_{t}\|^{2}]}{n\delta}}\right)\leq\frac{\mathrm{Var}[\|\epsilon_{t}\|^{2}]}{n\cdot\frac{\,\mathrm{Var}[\|\epsilon_{t}\|^{2}]}{n\delta}}=\delta,$ (A.5)
@@
$\leq(\kappa_{1}+\sqrt{\frac{\kappa_{2}}{n\delta I}})I\leq(\kappa_{1}+\sqrt{\kappa_{2}})I.$ (A.6)
The last inequality is because we can set $\delta$ such that $\delta nI<1$. Plug into Equation A.3, we have with probability higher than $1-\delta$, we obtain Equation A.1 and the lemma follows.
@@
Lemma A.2.
Under Assumption 5.3, for any $n$, there exists an event $A_{n}$ which occurs with probability at least $1-\sqrt{\frac{\log(16n)}{n}}$, such that, conditioning on $A_{n}$,
@@
The proof follows Lemma 1 in [39] that utilizes Dvoretzky-Kiefer-Wolfowitz inequality in [18]. ∎
Lemma A.3.
Under Assumption 5.1 5.3,with high probability,
@@
[…]
Using Lemma A.1 we have that with probability $1-\delta$,
$\sum_{t=T-n+1}^{T}|s_{t}-\hat{s}_{t}|$ $\leq n\omega.$ (A.7)
@@
$|S|\sqrt{\omega}\leq\sum_{t=T-n+1}^{T}|s_{t}-\hat{s}_{t}|\leq n\omega.$ (A.8)
@@
$\leq(2L+1)\sqrt{\omega}+2\sup_{x}\left|\widetilde{F}_{n+1}(x)-F_{s}(x)\right|,$ (A.9)
@@
$\left|\mathbb{P}\left(Y_{T+1}\in\mathcal{C}_{T+1}^{\alpha}\mid X_{T+1}=x_{T+1}\right)-(1-\alpha)\right|$ (A.10)
@@
$\leq$ $\mathbb{P}\left(|F_{s}(s_{T+1})-(1-\alpha)|\leq|\widehat{F}_{n+1}(\widehat{s}_{T+1})-F_{s}(s_{T+1})|\right).$ (A.11)
Based on Lemma A.2, we can define the event $A_{n},\mathbb{P}(A_{n})\geq 1-\frac{\log(16n)}{n}$, conditional on $A_{n}$, we have:
$\sup_{x}\left|\widetilde{F}_{n+1}(x)-F_{s}(x)\right|\leq\sqrt{\frac{\log(16n)}{n}},$ (A.12)
Hence, we can write Equation A.11 as
@@
$\leq$ $\mathbb{P}(|F_{s}(s_{T+1})-(1-\alpha)|\leq|\widehat{F}_{n+1}(\widehat{s}_{T+1})-F_{s}(\widehat{s}_{T+1})|+|F_{s}(\widehat{s}_{T+1})-F_{s}(s_{T+1})||A_{n})+\frac{\log(16n)}{n}.$ (A.13)
@@
$\leq$ $(2L+1)\sqrt{\omega}+3\sqrt{\frac{\log(16n)}{n}}+L\omega.$ (A.14)
The last equation exists because of Lemma A.3 A.1 and Equation A.12.
@@
$\leq$ $(4L+2)\sqrt{\omega}+6\sqrt{\frac{\log(16n)}{n}}+2L\omega.$ (A.15)
Plug into Equation A.11, we have
@@
$\leq$ $(4L+2)\sqrt{\omega}+6\sqrt{\frac{\log(16n)}{n}}+2L\omega+\frac{\log(16n)}{n}.$ (A.16)
A.2 Proof of Theorem 5.5
@@
Since $\epsilon\sim N(0,A_{*}^{-1})$ and $A\succ 0$ by Cholesky decomposition, we can write $A_{*}^{-1}=LL^{\top}$ where $L$ is a lower triangular matrix. Define the matrix $B=L^{\top}AL$, we can rewrite the  Equation 5.3 as
$\min_{B\succ 0,\|B\|\leq 1}Q_{1-\alpha}^{I/2}(x^{\top}Bx)\det(L)[\det(B)]^{-1/2},$ (A.17)
@@
To further solve the optimization problem, we look at the eigenvalue of $B$, suppose $B=O\text{diag}(\lambda_{1},\lambda_{2},\cdots,\lambda_{d})O^{\top}$ and $O$ is an orthogonal matrix.
$\min_{\lambda_{i}>0,\max_{1\leq i\leq I}\lambda_{i}\leq 1}Q^{I}_{1-\alpha}\left(\sum_{i=1}^{I}\lambda_{i}x_{i}^{2}\right)\prod_{i=1}^{d}\lambda_{i}^{-\frac{1}{2}},$ (A.18)
and $\{x_{i}\}_{1\leq i\leq d}$ are i.i.d. random variables $x_{i}\sim N(0,1)$.
Now we would like to prove that the above optimization problem is solved when $\lambda_{1}=\lambda_{2}\cdots=\lambda_{d}=1$ Consider the following “pairwise smoothing” step: for a pair $(\lambda_{i},\lambda_{j}),0<\lambda_{i}<\lambda_{j}\leq 1$, let $\delta>0$ be small consider the following new eigenvalue set
$\lambda_{k}^{\prime}:=\begin{cases}\lambda_{k}&k\neq i,j\\ \lambda_{i}+\delta&k=i\\ \lambda_{j}-\delta&k=j.\end{cases}$
Note that here we make $\delta$ small such that $\max_{1\leq i\leq d}\lambda^{\prime}_{i}\leq 1$. Since $\lambda_{i}^{\prime}\lambda_{j}^{\prime}>\lambda_{i}\lambda_{j}$, so $\prod_{i=1}^{d}\lambda_{i}^{\prime-\frac{1}{2}}<\prod_{i=1}^{d}\lambda_{i}^{-\frac{1}{2}}$. On the other hand, by rearrangement inequality and majorization theorem [22], the “pairwise smoothing” step strictly decreases or at least does not increase the tail quantile function $Q_{1-\alpha}$. By applying the pairwise smoothing repeatedly, we can show that Equation A.18 is minimized when $\lambda_{1}^{\prime}=\lambda_{2}^{\prime}=\cdots=\lambda_{I}^{\prime}=\frac{\sum_{i}{\lambda_{i}}}{I}$. Hence $B=I_{d}$ can solve the optimization problem Equation A.17, subsequently $A=(L^{\top})^{-1}L^{-1}=A_{*}$. ∎
A.3 Tailup model
Lemma A.4.

### v2

@@ 6.2 Real Data Study
Table 1 clearly demonstrates that our method consistently outperforms all baseline CP techniques across diverse backbone models and road topologies, achieving significantly higher coverage and superior efficiency. Further comparisons against CopulaTSCP and DeepSTUQ are detailed in Appendix D.5, as those methods are less fitted for our problem setting.
Our method exhibits robustness to the choice of the hyperparameter $\lambda$. In the first two subplots of Figure 3, setting $\lambda=0$ reduces our algorithm to sole usage of sample covariance matrix, equivalent to our strongest baseline, MultiDimSPCI. Our coverage is greater than MultiDimSPCI with arbitrary hyperparameter $\lambda$. Further, across all calibration sample sizes $n$, selecting any $\lambda\in[0.3,0.9]$ consistently yields both higher coverage and greater(lower) efficiency. This underscores the critical contribution of topological information and demonstrates STACI’s insensitivity to $\lambda$.
The incorporation of topology-induced covariance matrix is pivotal even without ACI ($\gamma=0$), as shown in the last two subplots in Figure 3. STACI can obviously lift coverage from under 8787% to surpass the desired 9595% level with only a marginal efficiency cost. This indicates that under inherent temporal covariance shifts, utilizing topological information offers a robust remedy to under-coverage problem while maintaining highly informative predictions. Moreover, as demonstrated in Section D.4 in the Appendix, STACI can also outperform in an offline setting, where the covariance matrix is estimated once and then held fixed at the beginning of test time. We discuss the scalability of STACI in term of time length and graph size in Section 7 and in Appendix D.7, respectively.
@@
In conclusion, the estimate of the covariance matrix can benefit from both topology and samples, compared with relying on any single resource. On one hand, with limited finite calibration sample $n$, the topology-based estimator offers a more stable structure as it possess fewer parameters. It can alleviate the temporal distribution shift and the resulting under-cover problem, consistent with our theoretical analysis. The sample covariance, on the other hand, captures the actual spatial patterns from samples in the calibration data and gives higher efficiency, but it can lose coverage guarantee if the distribution changes. By blending these two estimates and adaptively adjusting the significance level, STACI can effectively maintain desired coverage and smaller volumes.
@@
Figure 3: Comparison of Coverage and Efficiency for PeMS data with different belief weight $\lambda$ and calibration set size $n$, with adaptive step size $\gamma=0.01$ (upper) and 0 (lower). The pre-determined coverage threshold of 9595% is shown by a horizontal gray dotted line.
7 Limitation
STACI targets joint uncertainty quantification on moderate-size subgraphs ($2\sim 30$D). This choice is deliberate: performing UQ in very high-dimensional output spaces (hundreds of coordinates or more) is statistically ill-posed. Directly constructing geometric prediction sets in the full output space is vulnerable to the curse of dimensionality verleysen2005curse. In particular, hyper-rectangles formed by marginal intervals can inflate to near-vacuous volumes even at low dimensions as in Table 1.
In high-dimensional regimes, common practice is to avoid explicit geometric sets and instead use generative or sampling-based UQ le2016sampling; bohm2019uncertainty. For instance, image UQ, a prototypical high-dimensional setting, typically samples plausible outcomes from a learned posterior ekmekci2023quantifying rather than delineating a set in pixel space. In the same spirit, STACI focuses on subgraphs of manageable joint dimensionality so that (i) the resulting joint sets remain informative, (ii) the coverage–efficiency trade-off is non-degenerate. This design also matches how practitioners assess risk: traffic engineers often analyze corridors with 10–30 sensors rather than entire city-wide networks, and hydrologists frequently study clusters of $\sim$20 gauges when evaluating localized flood risk.
@@ 8 Conclusion
We proposed STACI, an adaptive conformal prediction framework for stream networks. Theoretically, we established coverage guarantees and demonstrated the model’s ability to minimize inefficiency under mild conditions. Empirically, STACI produced smaller prediction sets while maintaining valid coverage across both (stationary) simulated data and (non-stationary) real-world traffic data.
Future work includes three potential directions. (i) STACI can be extended to general spatio-temporal graphs by replacing $\Sigma_{\mathcal{G}}$ with alternative network parameterizations, enabling the development of novel methods that effectively exploit topological structures in broader spatio-temporal settings; (ii) stronger theoretical guarantees such as finite-sample coverage bounds when adaptively calibrating the significance level, could be developed by imposing assumptions on error distribution shifts (e.g., first-order differencing stationarity); and (iii) the present formulation does not aim to provide full-graph joint sets over hundreds of nodes. One might build corridor-level joint sets and composing them with rigorous controls on cross-correlation to scale coverage beyond the subgraph level.
Acknowledgements
We thank Aravindan Vijayaraghavan, Shuwen Chai, Yifan Wu for helpful discussions. We also thank anonymous reviewers for constructive feedback.
Contents
Appendix A Theoretical proofs
A.1 Proof of Theorem 1
Theorem 1(validity). Under the assumptions stated above, the proposed method satisfies the following conditional coverage guarantee:
$\left|\mathbb{P}(Y_{T+1}\in\mathcal{C}_{T+1}(\alpha)|X_{T+1}=x)-(1-\alpha)\right|\leq(4L+2L\sqrt{\omega}+2)\sqrt{\omega}+6\sqrt{\frac{\log(16n)}{n}}+\frac{\log(16n)}{n},$
where
@@ A.1 Proof of Theorem 1
[…]
@@
$\leq$ $r\|\Delta\|^{2}+2r\|\Delta\|\|\epsilon\|+o(g(n))\|\epsilon\|^{2}.$ (9)
The inequality (i) exists because of the Cauchy-schwartz inequality and Assumption 2. Hence, by Assumption 1, we have
@@
$\leq$ $rn\nu_{n}^{2}+2r\sqrt{n\nu_{n}^{2}(\sum_{t=T-n+1}^{T}\|\epsilon_{t}\|^{2})}+o(g(n))\sum_{t=T-n+1}^{T}\|\epsilon_{t}\|^{2}.$ (10)
From Assumption 3, we have
$\mathbb{E}\left[\frac{1}{n}\sum_{t=T-n+1}^{T}\|\epsilon_{t}\|^{2}\right]=\frac{1}{n}\sum_{t=T-n+1}^{T}\mathbb{E}[\|\epsilon_{t}\|^{2}]\leq\kappa_{1}I.$ (11)
@@
$\mathbb{P}\left(\frac{1}{n}\sum_{t=T-n+1}^{T}\|\epsilon_{t}\|^{2}-\mathbb{E}[\|\epsilon_{t}\|^{2}]\geq\sqrt{\frac{\,\mathrm{Var}[\|\epsilon_{t}\|^{2}]}{n\delta}}\right)\leq\frac{\mathrm{Var}[\|\epsilon_{t}\|^{2}]}{n\cdot\frac{\,\mathrm{Var}[\|\epsilon_{t}\|^{2}]}{n\delta}}=\delta,$ (12)
@@
$\leq(\kappa_{1}+\sqrt{\frac{\kappa_{2}}{n\delta I}})I\leq(\kappa_{1}+\sqrt{\kappa_{2}})I.$ (13)
The last inequality is because we can set $\delta$ such that $\delta nI<1$. Plug into eq.˜10, we have with probability higher than $1-\delta$, we obtain eq.˜8 and the lemma follows.
@@
Lemma 2.
Under Assumption 3, for any $n$, there exists an event $A_{n}$ which occurs with probability at least $1-\sqrt{\frac{\log(16n)}{n}}$, such that, conditioning on $A_{n}$,
@@
The proof follows Lemma 1 in [xu2023conformal] that utilizes Dvoretzky-Kiefer-Wolfowitz inequality in [kosorok2008introduction]. ∎
Lemma 3.
Under Assumption 1, Assumption 3,with high probability,
@@
The proof is similar to Lemma $B.6$ in [xuconformal], and is written here for completeness.
Using Lemma 1 we have that with probability $1-\delta$,
$\sum_{t=T-n+1}^{T}|s_{t}-\hat{s}_{t}|$ $\leq n\omega.$ (14)
@@
$|S|\sqrt{\omega}\leq\sum_{t=T-n+1}^{T}|s_{t}-\hat{s}_{t}|\leq n\omega.$ (15)
@@
$\leq(2L+1)\sqrt{\omega}+2\sup_{x}\left|\widetilde{F}_{n+1}(x)-F_{s}(x)\right|,$ (16)
@@
$\left|\mathbb{P}\left(Y_{T+1}\in\mathcal{C}_{T+1}^{\alpha}\mid X_{T+1}=x_{T+1}\right)-(1-\alpha)\right|$ (17)
@@
$\leq$ $\mathbb{P}\left(|F_{s}(s_{T+1})-(1-\alpha)|\leq|\widehat{F}_{n+1}(\widehat{s}_{T+1})-F_{s}(s_{T+1})|\right).$ (18)
Based on Lemma 2, we can define the event $A_{n},\mathbb{P}(A_{n})\geq 1-\frac{\log(16n)}{n}$, conditional on $A_{n}$, we have:
$\sup_{x}\left|\widetilde{F}_{n+1}(x)-F_{s}(x)\right|\leq\sqrt{\frac{\log(16n)}{n}},$ (19)
Hence, we can write eq.˜18 as
@@
$\leq$ $\mathbb{P}(|F_{s}(s_{T+1})-(1-\alpha)|\leq|\widehat{F}_{n+1}(\widehat{s}_{T+1})-F_{s}(\widehat{s}_{T+1})|+|F_{s}(\widehat{s}_{T+1})-F_{s}(s_{T+1})||A_{n})+\frac{\log(16n)}{n}.$ (20)
@@
$\leq$ $(2L+1)\sqrt{\omega}+3\sqrt{\frac{\log(16n)}{n}}+L\omega.$ (21)
The last equation exists because of Lemma 3 1 and eq.˜19.
@@
$\leq$ $(4L+2)\sqrt{\omega}+6\sqrt{\frac{\log(16n)}{n}}+2L\omega.$ (22)
Plug into eq.˜18, we have
@@
$\leq$ $(4L+2)\sqrt{\omega}+6\sqrt{\frac{\log(16n)}{n}}+2L\omega+\frac{\log(16n)}{n}.$ (23)
A.2 Proof of Theorem 2
Theorem 2 (Efficiency). There exists $0<\alpha_{0}<1$, such that when $\alpha<\alpha_{0}$, the optimal solution to the minimization problem is given by:
$A_{*}\coloneqq\arg\min_{A\succ 0}V(A,Q_{1-\alpha}(\epsilon^{\top}A\epsilon)),$ (24)
where $\epsilon\sim\mathcal{N}(0,A_{*}^{-1})$.
Lemma 4 (Uniform tail threshold via Dirichlet–Jensen).
Let $S_{\lambda}=\sum_{i=1}^{I}\lambda_{i}Z_{i}^{2}$ with $Z_{i}\overset{i.i.d}{\sim}\mathcal{N}(0,1)$ and $\lambda\in\Delta^{I-1}$. Then $S_{\lambda}\overset{d}{=}T\,U_{\lambda}$ with $T\sim\chi^{2}_{I}$ independent of $U_{\lambda}=\sum_{i}\lambda_{i}V_{i}$, where $V\sim\mathrm{Dirichlet}(\tfrac{1}{2},\ldots,\tfrac{1}{2})$. For all $\lambda$ we have $0<U_{\lambda}\leq 1$ and $\mathbb{E}[U_{\lambda}]=1/I$. Moreover, for $s\geq I+2$, the map $u\mapsto\bar{F}_{I}(s/u)$ is convex on $(0,1]$ (where $\bar{F}_{I}$ is the $\chi^{2}_{I}$ survival function). Consequently,
$\mathbb{P}(S_{\lambda}\geq s)=\mathbb{E}\big[\bar{F}_{I}(s/U_{\lambda})\big]\geq\bar{F}_{I}\!\big(s/\mathbb{E}[U_{\lambda}]\big)=\bar{F}_{I}(Is)=\mathbb{P}(S_{\lambda^{\star}}\geq s).$
Equivalently, with $p^{*}(I):=F_{\chi^{2}_{I}}\big(I(I+2)\big)\in(0,1)$, we have
$Q_{p}(S_{\lambda^{\star}})\ \leq\ Q_{p}(S_{\lambda}),\qquad\forall\,\lambda\in\Delta^{I-1},\ \forall\,p>p^{*}(I).$
@@ A.2 Proof of Theorem 2
[…]
Proof of Theorem 2
Proof.
By the definition of ellipsoid volume, we have
$V(A,Q_{1-\alpha}(\epsilon^{\top}A\epsilon))=Constant\times Q_{1-\alpha}(\epsilon^{\top}A\epsilon)^{I/2}[\det(A)]^{-1/2}$ (25)
Since the optimization problem is invariant to the rescaling of the scalar (that is, $A$ and any positive scalar multiple $cA$, where $c>0$, produce the same mathematical solution), additional constraints, such as the bounding of the matrix norm $A\leq 1$, can be imposed without loss of generality. Note that $\epsilon\sim N(0,A_{*}^{-1})$ and $A\succ 0$ by Cholesky decomposition, we can write $A_{*}^{-1}=LL^{\top}$ where $L$ is a lower triangular matrix. Define the matrix $B=L^{\top}AL$, we can rewrite the  eq.˜25 as
$\min_{B\succ 0}Q_{1-\alpha}^{I/2}(x^{\top}Bx)\det(L)[\det(B)]^{-1/2},$ (26)
@@
To further solve the optimization problem, we look at the eigenvalue of $B$, suppose $B=O\text{diag}(\lambda_{1},\lambda_{2},\cdots,\lambda_{I})O^{\top}$ and $O$ is an orthogonal matrix. Since the optimization value is invariant to different scaling of $B$, we are imposing additional constrains on $\lambda_{i}$.
$\min_{\lambda_{i}\geq 0,\sum_{i=1}^{I}\lambda_{i}=1}Q^{I/2}_{1-\alpha}\left(\sum_{i=1}^{I}\lambda_{i}x_{i}^{2}\right)\prod_{i=1}^{I}\lambda_{i}^{-\frac{1}{2}},$ (27)
and $\{x_{i}\}_{1\leq i\leq I}$ are i.i.d. random variables $x_{i}\sim N(0,1)$.
Now we would like to prove that the above optimization problem is solved when $\lambda^{*}=(\frac{1}{I},\cdots,\frac{1}{I})$. Note that by Cauchy-Schwartz Inequality $\prod_{i=1}^{I}\lambda_{i}^{-\frac{1}{2}}$ is minimized when $\lambda_{1}=\lambda_{2}\cdots=\lambda_{d}=\frac{1}{I}$, by Lemma 4, $Q_{1-\alpha}(\sum\lambda_{i}x_{i}^{2})$ is minimized in $\lambda^{*}$, for $\alpha<\alpha_{0}$ and $\alpha_{0}$ is decided according to $I$. Theorem 2 follows.
A.3 Tail-up model
Lemma 5.